OpenStudy (anonymous):

The perimeter of a rectangle is 42 ft. The length is 5 ft. longer than the width. Let x be the length and y be the width. 2x+2y=? x=y+? What is the length? What is the width?

OpenStudy (anonymous):

alright, so one way to write the perimeter is 2 times the length of one side plus 2 times the length of the next side for a rectangle, right? So 2x+2y=42

OpenStudy (anonymous):

if y is the width and we know that the length (x) is the width plus 5, then x=y+5, right?

OpenStudy (anonymous):

because x=y+5, you can take y+5 and put it where x is in the first equation and then solve that one for y, which is the width, remember.

OpenStudy (anonymous):

and because the length is y+5, add 5 to what you got for y and you have the length!

OpenStudy (anonymous):

Thank You very MUCH!